OpenStudy (anonymous):

2. What is the actual straight-line distance between two points if one lies 20 miles west and 16 miles north of the other? What angle does this line make with the north? Please show steps to get the right answer.

OpenStudy (anonymous):

Have you drawn a triangle on a piece of paper yet? Are you familiar with the Pythagorean theorem and the "soh cah toa" rule of thumb?

OpenStudy (anonymous):

Yes, I am aware of trignometric method. I got the answer for the first question, but I'm not quite sure about how to get the answer for second question.

OpenStudy (anonymous):

How can I use those rules to find the answer for second question?

OpenStudy (anonymous):

considering a central reference point.......... this will make a right angled triangle with sides 20miles 16miles and the hypotennuse which makes the distance between the 2 points , then actual straight line distance=\[\sqrt{(20^{2}+16^{2}}\] = distance in miles tan x=oppositeside/adjuscent side=20/16=5/4 angle x=tan inverse 5/4= answer

OpenStudy (anonymous):

Wow. Thank you fellow owl.

OpenStudy (anonymous):

Actually, I think the second answer is 90 - tan inverse 5/4. Isn't it?

OpenStudy (anonymous):

sorry the opposite side is 16miles and adjuscent side is 20 miles...... anyway U got it ......sorry ....my mistake was that Itook the sides wrongly......... U understood the way Right.